# On Computational Limits and Provably Efficient Criteria of Visual Autoregressive Models: A Fine-Grained Complexity Analysis

### v1

@@ 5 Provably Efficient Criteria
Section 5.1 details the running time of the fast $\operatorname{\mathsf{VAR}}$ Transformer, feature map reconstruction block, and Fast VQ-VAE Decoder. In Section 5.2, we analyze the error propagation in both the Fast $\operatorname{\mathsf{VAR}}$ Transformer and the Fast VQ-VAE Decoder. Section 5.3 presents our findings regarding the existence of an almost quadratic time algorithm.
@@ 5.1 Running Time
Lemma 5.1 (Running time of Fast $\operatorname{\mathsf{VAR}}$ Transformer, informal version of Lemma E.3).
Assuming the conditions below are satisfied:
  • Let $K$ denote the total number of the generated token maps.
  • Let $k\in[K]$.
  • Let $r_{1}\in\mathbb{R}^{1\times 1\times d}$ denote the first scale token map.
  • Let $d=O(\log(n))$.
@@
Lemma 5.2 (Running time of Feature Map Reconstruction Layer, informal version of Lemma E.4).
Assuming the conditions below are satisfied:
  • Let $K$ denote the total number of generated token maps.
  • Let $k\in[K]$.
  • Let $r_{1}\in\mathbb{R}^{1\times 1\times d}$ denote the first scale token map.
  • Let $d=O(\log(n))$.
@@
Lemma 5.3 (Running time of Fast VQ-VAE Decoder, informal version of Lemma E.6).
Assuming the conditions below are satisfied:
  • Define $k_{1},k_{2},k_{3}\in\mathbb{N}$ as constant numbers.
  • Let $X\in\mathbb{R}^{n\times n\times d}$ represent the input feature map.
  • We assume there are $k_{1}$ up-interpolation layers $\phi_{\rm up}$ defined in Definition 3.2.
  • Given a feature map $M\in\mathbb{R}^{h\times w\times d}$. For $i\in[k_{1}]$, we assume $i$-th up-interpolation layer’s output $\phi^{i}_{\rm up}(M)\in\mathbb{R}^{h\times w\times O(d)}$.
  • Let $d=O(\log(n))$.
@@ 5.2 Error Propagation Analysis
Here, we present the error analysis results of the fast Transformer layer and the fast VQ-VAE Decoder.
We begin with analyzing the error propagation of $\mathsf{AAttC}$ (see Definition 4.2).
Lemma 5.4 (Error analysis of $\mathsf{AAttC}(X^{\prime})$ and $\mathsf{AAttC}(X)$, informal version of Lemma C.3).
Assuming the conditions below are satisfied:
  • Define $X\in\mathbb{R}^{n\times d}$ as the input matrix.
  • Let $\epsilon\in(0,0.1)$ denote the approximation error.
  • Suppose we have $\|X^{\prime}-X\|_{\infty}\leq\epsilon$.
  • Let $M>1$.
  • Assume the value of each entry in matrices can be bounded by $M$.
  • Let the polynomial approximation of attention matrix $\mathsf{AAttC}(X)$ as given in Definition 4.2.
  • Let $f$ be a polynomial with degree $g$.
Then, we can show that
$\|\mathsf{AAttC}(X^{\prime})-\mathsf{AAttC}(X)\|_{\infty}\leq O(kM^{g+2}d)\cdot\epsilon$
Then, we move on to analyzing the approximation error between $\mathsf{AAttC}(X^{\prime})$ and the ground truth $\mathsf{Attn(X)}$, where $X^{\prime}$ denotes the approximated version of the original input $X$.
Lemma 5.5 (Error analysis of $\mathsf{AAttC}(X^{\prime})$ and $\mathsf{Attn}(X)$, informal version of Lemma C.4).
Assuming the conditions below are satisfied:
  • Define $X\in\mathbb{R}^{n\times d}$ as the input matrix.
  • Let $\epsilon\in(0,0.1)$ denote the approximation error.
  • Suppose we have $\|X^{\prime}-X\|_{\infty}\leq\epsilon$.
  • Let $M>1$.
  • Assume the value of each entry in matrices can be bounded by $M$.
  • Let the attention matrix $\mathsf{Attn}(X)\in\mathbb{R}^{n\times n}$ as given in Definition 3.3.
  • Let the polynomial approximation of attention matrix $\mathsf{AAttC}(X)\in\mathbb{R}^{n\times n}$ as given in Definition 4.2.
@@
Then, we can show that
$\|\mathsf{AAttC}(X^{\prime})-\mathsf{Attn}(X)\|_{\infty}\leq O(kM^{g+1}d)\cdot\epsilon$
With the above foundations, we are ready to move to analyzing the approximation error of one fast $\operatorname{\mathsf{VAR}}$ layer (see Definition 3.6).
Lemma 5.6 (Error propagation analysis for one $\operatorname{\mathsf{VAR}}$ Transformer Layer, informal version of Lemma C.7).
Assuming the conditions below are satisfied:
  • Define $X\in\mathbb{R}^{n\times d}$ as the input data matrix.
  • Let $\epsilon\in(0,0.1)$ denote the approximation error.
  • Suppose we have $\|X^{\prime}-X\|_{\infty}\leq\epsilon$.
  • Let the $\operatorname{\mathsf{VAR}}$ Transformer Layer $\mathsf{F}_{\mathrm{var}}$ as given in Definition 3.5.
  • Let the Fast $\operatorname{\mathsf{VAR}}$ Transformer Layer $\mathsf{F}_{\rm fvar}$ as given in Definition 3.6.
  • Let $M>1$.
  • Assume the value of each entry in matrices can be bounded by $M$.
  • Let the attention matrix $\mathsf{Attn}(X)\in\mathbb{R}^{n\times n}$ be defined as Definition 3.3.
  • Let the polynomial approximation of attention matrix $\mathsf{AAttC}(X)\in\mathbb{R}^{n\times n}$ be defined as Definition 4.2.
  • Let $f$ be a polynomial with degree $g$.
Then, we can show that
$\|\mathsf{F}_{\rm fvar}(X^{\prime})-\mathsf{F}_{\mathrm{var}}(X)\|_{\infty}\leq O(kM^{g+1}d)\cdot\epsilon$
Then, we show that the $\operatorname{\mathsf{VAR}}$ Transformer will introduce additive error bounded by $1/\operatorname{poly}(n)$.
Lemma 5.7 (Error analysis for the Fast $\operatorname{\mathsf{VAR}}$ Transformer, informal version of Lemma C.8).
Assuming the conditions below are satisfied:
  • Define $X_{0}\in\mathbb{R}^{1\times d}$ as the initial input data matrix.
  • Define $r\in\mathcal{N}$ as the total iteration of the Fast $\operatorname{\mathsf{VAR}}$ model.
  • For $i\in[r]$, let $T_{i}(X_{0})$ denote the intermediate variable of the $\operatorname{\mathsf{VAR}}$ model, as defined in Definition 3.7.
  • For $i\in[r]$, let $\widetilde{T}_{i}(X_{0})$ denote the intermediate variable of the fast $\operatorname{\mathsf{VAR}}$ model, as defined in Definition 3.8.
  • Assume $T_{r}(X_{0})\in\mathbb{R}^{O(n^{2})\times d}$ is final output of the $\operatorname{\mathsf{VAR}}$ Transformer.
  • Assume $\widetilde{T}_{r}(X_{0})\in\mathbb{R}^{O(n^{2})\times d}$ is final output of the Fast $\operatorname{\mathsf{VAR}}$ Transformer.
  • Assume each element of the matrices is encoded in $O(\log n)$ bits.
Then, we can show that the error bound of the final output $\widetilde{T}_{r}(X_{0})$ as
$\|\widetilde{T}_{r}(X_{0})-T_{r}(X_{0})\|_{\infty}\leq 1/\operatorname{poly}(n)$
@@ 5.3 Existence of Almost Quadratic Time Algorithm
By combining the result of Lemma 5.1, Lemma 5.2, Lemma 5.3 and Lemma 5.7, we can easily derive the proof. ∎
@@ 6 Discussion
This work provides a foundational framework for understanding and overcoming the computational bottlenecks in generative models. It will serve as a key resource for researchers striving to design the next generation of efficient autoregressive models. By addressing the limitations of current models and offering clear guidance on how to optimize them, we hope to inspire more efficient and scalable solutions for a wide array of machine learning applications, extending far beyond visual generation.
@@ 7 Conclusion
Roadmap. Section A introduces key notations. Section B provides an analysis of the perturbation error in different VAR model modules. Section C details the error analysis for the VAR Transformer. In Section D, we present the error analysis of the VQVAE decoder. In Section E, we evaluate the running time of VAR models and fast VAR models.
@@ Appendix A Notations
Appendix B Perturbation Error Analysis
Section B.1 focuses on the perturbation error of linear interpolation. In Section B.2, we explore the error incurred when raising an approximated value to a power. Section B.3 analyzes the perturbation error of approximating the product of two values, while Section B.4 addresses the error in approximating the product of $d$ values.
B.1 Perturbation Error for Linear Interpolation
First, we state a fact that demonstrates that the simple linear interpolation operation does not introduce additional error.
Fact B.1.
Assuming the conditions below are satisfied:
  • Let $a,b\in\mathbb{R}$.
  • Define $a^{\prime},b^{\prime}\in\mathbb{R}$ as the approximated value of $a,b$.
  • Let $|a|\leq R$ and $|a^{\prime}|\leq R$.
  • Let $|b|\leq R$ and $|b^{\prime}|\leq R$.
  • Let $\delta\in(0,0.1)$ denote the approximation error.
  • Suppose $|a-a^{\prime}|\leq\delta$ and $|b-b^{\prime}|\leq\delta$.
Then we have, for any $w\in(0,1)$
$|(wa^{\prime}+(1-w)b^{\prime})-(wa+(1-w)b)|\leq\delta$
Proof.
We have
$~{}|(wa^{\prime}+(1-w)b^{\prime})-(wa+(1-w)b)|$
$=$ $~{}|w(a^{\prime}-a)+(1-w)(b^{\prime}-b)|$
$\leq$ $~{}w|a^{\prime}-a|+(1-w)|b^{\prime}-b|$
$\leq$ $~{}w\cdot\delta+(1-w)\cdot\delta$
$=$ $~{}\delta$
We derive the first step using basic algebra, the second step using the triangle inequality and the fact that $w\in(0,1)$, the third step using the conditions of the lemma, and the last step using basic algebra.
B.2 Perturbation Error of Power Function
We then present several lemmas that will be utilized in the subsequent proof.
Fact B.2.
Assuming the conditions below are satisfied:
  • Let $a\in\mathbb{R}$.
  • Let $a^{\prime}\in\mathbb{R}$ denote the approximation value of $a$.
  • Let $|a|\leq R$.
  • Let $|a^{\prime}|\leq R$.
  • Define $\delta\in(0,0.1)$ as the approximation error.
  • Suppose $\delta\leq|a|$.
  • Suppose $|a^{\prime}-a|\leq\delta$.
  • Let $m\in\mathbb{N}$.
  • Let $C:=(2^{m}-1)\cdot R^{m-1}$.
Then we have
$|(a^{\prime})^{m}-a^{m}|\leq C\cdot\delta$
Proof.
Case 1: We firstly consider the case $|(a^{\prime})^{m}-a^{m}|\leq|(a+\delta)^{m}-a^{m}|$. Then we have:
$|(a^{\prime})^{m}-a^{m}|\leq$ $~{}|(a+\delta)^{m}-a^{m}|$
$\leq$ $~{}|\sum_{i=0}^{m}\binom{m}{i}a^{m-i}\delta^{i}-a^{m}|$
$=$ $~{}|\sum_{i=1}^{m}\binom{m}{i}a^{m-i}\delta^{i}|$
$\leq$ $~{}\sum_{i=1}^{m}|\binom{m}{i}a^{m-i}\delta^{i}|$
The first step above is based on the condition of this lemma, the second relies on binomial expansion, the third uses simple algebra, and the last step follows from the triangle inequality.
Then, we keep on showing that:
$\sum_{i=1}^{m}|\binom{m}{i}a^{m-i}\delta^{i}|=$ $~{}\sum_{i=1}^{m}\binom{m}{i}\cdot|a^{m-i}|\cdot\delta^{i}$
$\leq$ $~{}\sum_{i=1}^{m}\binom{m}{i}\cdot|a^{m-1}|\cdot\delta$
$\leq$ $~{}(2^{m}-1)\cdot R^{m-1}\cdot\delta$
$=$ $~{}C\cdot\delta$
The first step above is based on $\delta>0$ and $\binom{m}{i}>0$. The second step is derived from the condition that $\delta\leq|a|$. The third step is due to the fact that $\sum_{i=1}^{m}\binom{m}{i}=2^{m}-1$ and the condition that $|a|\leq R$. The final step is trivially from the definition of $C$.
Case 2: Then we consider the case $|(a^{\prime})^{m}-a^{m}|\leq|(a-\delta)^{m}-a^{m}|$. Then we have:
$|(a^{\prime})^{m}-a^{m}|\leq$ $~{}|(a-\delta)^{m}-a^{m}|$
$\leq$ $~{}|\sum_{i=0}^{m}\binom{m}{i}a^{m-i}(-\delta)^{i}-a^{m}|$
$=$ $~{}|\sum_{i=1}^{m}\binom{m}{i}a^{m-i}(-\delta)^{i}|$
$\leq$ $~{}\sum_{i=1}^{m}|\binom{m}{i}a^{m-i}(-\delta)^{i}|$
$=$ $~{}\sum_{i=1}^{m}|\binom{m}{i}a^{m-i}\delta^{i}|$
The first step relies on this lemma’s condition, the second on binomial expansion, the third on basic algebra, the fourth on triangle inequality, and the last on basic algebra.
Then, we can further show that:
$\sum_{i=1}^{m}|\binom{m}{i}a^{m-i}\delta^{i}|=$ $~{}\sum_{i=1}^{m}\binom{m}{i}\cdot|a^{m-i}|\cdot\delta^{i}$
$\leq$ $~{}\sum_{i=1}^{m}\binom{m}{i}\cdot|a^{m-1}|\cdot\delta$
$\leq$ $~{}(2^{m}-1)\cdot R^{m-1}\cdot\delta$
$=$ $~{}C\cdot\delta$
The first step above relies on the condition that $\delta>0$ and $\binom{m}{i}>0$. And the second step uses the condition that $\delta\leq|a|$. The third step derives from the fact that $\sum_{i=1}^{m}\binom{m}{i}=2^{m}-1$ and the condition that $|a|\leq R$, and the final step is from the definition of $C$.
B.3 Perturbation Error of Production of Two Terms
In this section, we introduce the perturbation error of the production of two terms.
Fact B.3.
Assuming the conditions below are satisfied:
  • Let $a,b\in\mathbb{R}$.
  • Define $a^{\prime},b^{\prime}\in\mathbb{R}$ as the approximated values of $a,b$.
  • Let $|a|\leq R$
  • Let $|a^{\prime}|\leq R$.
  • Let $|b|\leq R$
  • Let $|b^{\prime}|\leq R$.
  • Let $\delta\in(0,0.1)$ denote the approximation value.
  • Suppose we have $|a^{\prime}-a|\leq\delta$.
  • Suppose we have $|b^{\prime}-b|\leq\delta$.
Then we have
$|a^{\prime}b^{\prime}-ab|\leq 2R\cdot\delta$
Proof.
We have
$|a^{\prime}b^{\prime}-ab|=$ $~{}|(a^{\prime}b^{\prime}-a^{\prime}b)+(a^{\prime}b-ab)|$
$=$ $~{}|a^{\prime}(b^{\prime}-b)+b(a^{\prime}-a)|$
$\leq$ $~{}|a^{\prime}|\cdot\delta+|b|\cdot\delta$
$\leq$ $~{}2R\cdot\delta$
Starting with simple algebra for the first and second steps, the third step uses the triangle inequality along with $|a^{\prime}-a|\leq\delta$ and $|b-b^{\prime}|\leq\delta$, and the final step is based on $|a^{\prime}|\leq R$ and $b\leq R$. ∎
B.4 Perturbation Error of Production of $d$ Terms
In this section, we introduce the perturbation error of production of $d$ terms.
Fact B.4.
Assuming the conditions below are satisfied:
  • Let $a_{1},a_{2},\cdots,a_{d}\in\mathbb{R}$.
  • Define $a^{\prime}_{1},a^{\prime}_{2},\cdots,a^{\prime}_{d}\in\mathbb{R}$ as the approximation values of $a_{1},a_{2},\cdots,a_{d}$.
  • For every $i\in[d]$, we assume $|a_{i}|\leq R$ and $|a^{\prime}_{i}|\leq R$.
  • Let $\delta\in(0,0.1)$ denote the approximation error.
  • For every $i\in[d]$, we assume $|a_{i}-a^{\prime}_{i}|\leq\delta$.
Then we have
$|\prod_{i=1}^{d}a_{i}-\prod_{i=1}^{d}a^{\prime}_{i}|\leq d\cdot R^{d-1}\cdot\delta$
Proof.
The proof is carried out by mathematical induction as follows:
Consider the case $d=2$. We can show that
$|\prod_{i=1}^{2}a_{i}-\prod_{i=1}^{2}a^{\prime}_{i}|=$ $~{}|a_{1}a_{2}-a^{\prime}_{1}a^{\prime}_{2}|$
$\leq$ $~{}2R\cdot\delta$
The reasoning above follows from Fact B.3. Assume the statement is true for $d=k$:
$|\prod_{i=1}^{k}a_{i}-\prod_{i=1}^{k}a^{\prime}_{i}|\leq k\cdot R^{k-1}\cdot\delta$ (1)
consequently, we can demonstrate that:
$~{}|\prod_{i=1}^{k+1}a_{i}-\prod_{i=1}^{k+1}a^{\prime}_{i}|$
$=$ $~{}|(\prod_{i=1}^{k}a_{i})\cdot a_{k+1}-(\prod_{i=1}^{k}a^{\prime}_{i})\cdot a^{\prime}_{k+1}|$
$=$ $~{}|(\prod_{i=1}^{k}a_{i})\cdot(a_{k+1}-a^{\prime}_{k+1})+(\prod_{i=1}^{k}a_{i}-\prod_{i=1}^{k}a^{\prime}_{i})\cdot a^{\prime}_{k+1}|$
$\leq$ $~{}|(\prod_{i=1}^{k}a_{i})\cdot(a_{k+1}-a^{\prime}_{k+1})|+|(\prod_{i=1}^{k}a_{i}-\prod_{i=1}^{k}a^{\prime}_{i})\cdot a^{\prime}_{k+1}|$
$\leq$ $~{}R^{k}\cdot\delta+k\cdot R^{k-1}\cdot\delta\cdot R$
$=$ $~{}(k+1)\cdot R^{k}\cdot\delta$
Starting with simple algebra for the first and second steps, the third step follows from the triangle inequality, the fourth from the conditions of this lemma and Eq. (1), and the final step involves simple algebra.
Hence, we have
$|\prod_{i=1}^{d}a_{i}-\prod_{i=1}^{d}a^{\prime}_{i}|\leq d\cdot R^{d-1}\cdot\delta$
where this comes from mathematical induction. ∎
Appendix C Error Analysis of Visual Auto-Regressive Transformer
This section focuses on the error analysis of the $\operatorname{\mathsf{VAR}}$ model. In Section C.1, we introduce the Lipschitz property of a polynomial function. In Section C.2, we analyze the error propagation of inner product operation by giving two approximated inputs. In Section C.3, we analyze the error propagation of the fast attention $\mathsf{AAttC}(X)$. In Section C.4, we analyze the error between $\mathsf{AAttC}(X^{\prime})$ and $\mathsf{Attn}(X)$ where $X^{\prime}$ is the approximated value of $X$. In Section C.5, we conduct an error analysis of the up-interpolation layer. In Section C.6, we conduct the error analysis of the $\operatorname{\mathsf{VAR}}$ transformer. Finally, in Section C.7, we conduct the error analysis of the Fast $\operatorname{\mathsf{VAR}}$ transformer.
C.1 Lipschitz of Polynomial

### v2

@@ 4.2 Hardness of Approximate Attention Computation
Now, we are able to give the definition of Fa ast $\operatorname{\mathsf{VAR}}$ transformer.
Definition 4.3 (Fast $\operatorname{\mathsf{VAR}}$ transformer).
We define fast $\operatorname{\mathsf{VAR}}$ transformer as follows:
  • Assume the $\operatorname{\mathsf{VAR}}$ transformer has $m$ attention layers.
  • Let $\Phi_{\mathrm{up},r}$ denote the pyramid up-interpolation layer defined in Definition 3.3.
  • Let $\mathsf{AAttC}_{i}$ stand for the $i$-th approximate attention computation layer, which is defined in Definition 4.2.
  • Given an input token map $X_{\mathrm{init}}\in\mathbb{R}^{1\times d}$.
  • Let $l=\sum_{i=1}^{K}h_{i}w_{i}$
We define the fast $\operatorname{\mathsf{VAR}}$ transformer as the following
$\mathsf{FTF}(X_{\mathrm{init}}):=$ $~{}\mathsf{AAttC}_{K}\circ\Phi_{\mathrm{up},K-1}\circ\cdots\circ\mathsf{AAttC}_{2}\circ\Phi_{\mathrm{up},1}$
$\circ$ $~{}\mathsf{AAttC}_{1}(X_{\mathrm{init}})\in\mathbb{R}^{l\times d},$
In this expression, $\circ$ stands for functional composition.
@@ 4.3 Computational Limits of Fast $\operatorname{\mathsf{VAR}}$ Models
In Theorem 4.5, we show our hardness result. The theorem states that we can’t accurately approximate ($\epsilon=1/\operatorname{poly}(n)$) the output of $\operatorname{\mathsf{VAR}}$ model in $O(n^{4-\Omega(1)})$ time when $\mathsf{SETH}$ holds. This implies that, under the condition of $\mathsf{SETH}$, achieving an efficient approximation algorithm for the $\operatorname{\mathsf{VAR}}$ model within this time bound is infeasible. As a result, we set a lower bound on the complexity of approximating the output of the $\operatorname{\mathsf{VAR}}$ model, indicating that any algorithm aimed at approximating the model with a small error will need to operate in at least $O(n^{4-\Omega(1)})$ time in the worst case.
@@ 5 Provably Efficient Criteria
Section 5.1 details the running time of the fast $\operatorname{\mathsf{VAR}}$ Transformer, feature map reconstruction layer, and Fast VQ-VAE Decoder. In Section 5.2, we analyze the error propagation in both the Fast $\operatorname{\mathsf{VAR}}$ Transformer and the Fast VQ-VAE Decoder. Section 5.3 presents our findings regarding the existence of an almost quadratic time algorithm.
@@ 5.1 Running Time
Lemma 5.1 (Running time of Fast $\operatorname{\mathsf{VAR}}$ Transformer, informal version of Lemma D.3).
Assume the fast $\operatorname{\mathsf{VAR}}$ transformer defined in Definition 4.3 has $K$ attention layers. Let $k_{1}\in[K]$ Let $X_{\mathrm{init}}\in\mathbb{R}^{1\times 1\times d}$ denote the first scale token map. Let $\alpha>1$ denote the growth rate of the height and width of the token map at each level. Then for $k_{1}\in[K]$, the $k_{1}$-th token map $r_{k_{1}}\in\mathbb{R}^{\alpha^{k_{1}-1}\times\alpha^{k_{1}-1}\times d}$. Let $r_{K}\in\mathbb{R}^{n\times n\times d}$ denote the last scale token map, where $n=\alpha^{K-1}$. Let $d=O(\log n)$ denote the embedding size of each token.
@@
Proof.
Please refer to Lemma D.3 for the proof’s details. ∎
@@
Lemma 5.2 (Running time of Feature Map Reconstruction Layer, informal version of Lemma D.4).
Let $X_{\mathrm{init}}\in\mathbb{R}^{1\times 1\times d}$ denote the initial token map. Let $\alpha>1$ denote the growth rate of the height and width of the token map at each level. Then for $k\in[K]$, the $k$-th token map $r_{k}\in\mathbb{R}^{\alpha^{k-1}\times\alpha^{k-1}\times d}$. Let $r_{K}\in\mathbb{R}^{n\times n\times d}$ denote the last scale token map, where $n=\alpha^{K-1}$. Let $d=O(\log n)$ denote the embedding size of each token.
@@
Proof.
Please refer to Lemma D.4 for the proof’s details. ∎
@@
Lemma 5.3 (Running time of Fast VQ-VAE Decoder, informal version of Lemma D.6).
Let $k_{1},k_{2},k_{3}\in\mathbb{N}$ be constant numbers. Given $X\in\mathbb{R}^{n\times n\times d}$ as the input feature map. Let $d=O(\log n)$. Assume that there are $k_{1}$ up-interpolation layers $\phi_{\rm up}$ defined in Definition 3.2. Given a feature map $M\in\mathbb{R}^{h\times w\times d}$. For $i\in[k_{1}]$, we assume $i$-th up-interpolation layer’s output $\phi^{i}_{\rm up}(M)\in\mathbb{R}^{O(h)\times O(w)\times d}$. We assume there are $k_{2}$ approximate attention layers $\mathsf{AAttC}$ defined in Definition 4.2. Given a feature map $M\in\mathbb{R}^{h\times w\times d}$. For $i\in[k_{1}]$, the $i$-th approximate attention layer’s output $\mathsf{AAttC}(M)\in\mathbb{R}^{h\times w\times d}$. We assume there are $k_{3}$ convolution layers $\phi_{\rm conv}$ defined in Definition 3.6. Given a feature map $M\in\mathbb{R}^{h\times w\times d}$. For $i\in[k_{1}]$, we assume $i$-th convolution layer’s output $\phi^{i}_{\rm conv}(M)\in\mathbb{R}^{h\times w\times O(d)}$.
@@
Proof.
Please refer to Lemma D.6 for the proof’s details. ∎
@@ 5.2 Error Propagation Analysis
In this section, we present an error analysis introduced by the fast algorithm applied to the $\operatorname{\mathsf{VAR}}$ and VQVAE models.
Firstly, we can show that the model output error introduced by the fast algorithm for $\operatorname{\mathsf{VAR}}$ will not exceed $1/\operatorname{poly}(n)$.
Lemma 5.4 (Error analysis for the Fast $\operatorname{\mathsf{VAR}}$ Transformer, informal version of Lemma B.7).
If the following conditions hold:
  • Let $X_{\mathrm{init}}\in\mathbb{R}^{1\times d}$ denote the initial input token map.
  • Let $K\in\mathbb{N}$ represent the number of approximate attention layers in the $\operatorname{\mathsf{VAR}}$ model.
  • Let the $\operatorname{\mathsf{VAR}}$ transformer $\mathsf{TF}_{K}$ be defined as Definition 3.5.
  • Let the Fast $\operatorname{\mathsf{VAR}}$ Transformer $\mathsf{FTF}_{K}$ as given in Definition 4.3.
  • For $i\in[K]$, let $\mathsf{TF}_{i}(X_{\mathrm{init}})$ denote the output of the i-th iteration of the $\operatorname{\mathsf{VAR}}$ transformer.
  • For $i\in[K]$, let $\mathsf{FTF}_{i}(X_{\mathrm{init}})$ denote the output of the i-th iteration of the fast $\operatorname{\mathsf{VAR}}$ transformer.
  • Let $\mathsf{TF}_{K}(X_{\mathrm{init}})\in\mathbb{R}^{O(n^{2})\times d}$ denote the final output of the $\operatorname{\mathsf{VAR}}$ transformer.
  • Let $\mathsf{FTF}_{K}(X_{\mathrm{init}})\in\mathbb{R}^{O(n^{2})\times d}$ denote the final output of the fast $\operatorname{\mathsf{VAR}}$ transformer.
@@
Then, we can show that the error bound of the final output $\mathsf{FTF}_{K}(X_{\mathrm{init}})$ as
$\|\mathsf{FTF}_{K}(X_{\mathrm{init}})-\mathsf{TF}_{K}(X_{\mathrm{init}})\|_{\infty}\leq 1/\operatorname{poly}(n)$
Proof.
Please refer to Lemma B.7 for the proof’s details. ∎
Remark 5.5.
Since the modules in the Feature Map Reconstruction stage (Stage 2) only consist of an up-interpolation layer and a convolution layer, without any attention layers, the acceleration method proposed in this paper cannot be applied to this stage. Moreover, since the Feature Map Reconstruction phase only involves simple linear operations, it is trivial that this stage will not introduce more than $1/\operatorname{poly}(n)$ error.
We also present that the model output error introduced by the fast algorithm for the VQ-VAE Decoder will not exceed $1/\operatorname{poly}(n)$.
Lemma 5.6 (Error analysis of Fast VQ-VAE Decoder, informal version of Lemma C.2).
Let $X\in\mathbb{R}^{n\times d}$ denote the input matrix. Let the up-interpolation Layer $\phi_{\rm up}$ be defined as Definition 3.2. Let the convolution layer $\phi_{\rm conv}$ be defined as Definition 3.6. Let the attention layer $\mathsf{Attn}$ be defined as Definition 3.4. Let the fast attention layer $\mathsf{AAttC}$ be defined as Definition 4.2. Let the VQ-VAE Decoder be the composition of a constant number of up-interpolation layers, convolutions layers, and attention layers. Let the Fast VQ-VAE Decoder be defined as substituting all $\mathsf{Attn}$ layers in VQ-VAE with $\mathsf{AAttC}$ layers.
Proof.
Please refer to Lemma C.2 for the proof’s details. ∎
@@ 5.3 Existence of Almost Quadratic Time Algorithm
By combining the result of Lemma 5.1, Lemma 5.2, Lemma 5.3, Lemma 5.4 and Lemma 5.6, we can easily derive the proof. ∎
In Theorem 5.7, we show that we can accurately approximates ($\epsilon=1/\operatorname{poly}(n)$) the overall $\operatorname{\mathsf{VAR}}$ model output in almost quadratic time $n^{2+o(1)}$ under practical assumptions. By Lemma 5.1, Lemma 5.2 and Lemma 5.3, we know that the bottleneck in accelerating the runtime of the VAR model is the attention computation (The origin running time cost is $O(n^{4+o(1)})$). The insight of the theorem is that we apply approximate attention computation $\mathsf{AAttC}$ (The running time cost is $O(n^{2+o(1)})$) to replace the original attention computation $\mathsf{Attn}$. In this way, our methods solve the $\operatorname{\mathsf{VAR}}$ model acceleration. This result enables the $\operatorname{\mathsf{VAR}}$ model to significantly accelerate the inference process in image generation, making it more competitive in the field of image generation.
@@ 6 Discussion
This work provides a foundational framework for understanding and overcoming the computational bottlenecks in generative models. It will serve as a key resource for researchers striving to design the next generation of efficient autoregressive models. By addressing the limitations of current models and offering clear guidance on how to optimize them, we hope to inspire more efficient and scalable solutions for a wide array of machine-learning applications, extending far beyond visual generation.
@@ 7 Conclusion
Roadmap. Section A introduces key notations. Section B details the error analysis for the VAR Transformer. In Section C, we present the error analysis of the VQVAE decoder. In Section D, we evaluate the running time of VAR models and fast VAR models.
@@ Appendix A Notations
Appendix B Error Analysis of Fast Visual Auto-Regressive Transformer
This section focuses on the error analysis of the $\operatorname{\mathsf{VAR}}$ model. In Section B.1, we introduce the Lipschitz property of a polynomial function. In Section B.2, we analyze the error propagation of inner product operation by giving two approximated inputs. In Section B.3, we analyze the error propagation of the fast attention $\mathsf{AAttC}(X)$. In Section B.4, we analyze the error between $\mathsf{AAttC}(X^{\prime})$ and $\mathsf{Attn}(X)$ where $X^{\prime}$ is the approximated value of $X$. In Section B.5, we conduct an error analysis of the up-interpolation layer. In Section B.6, we conduct the error analysis of the $\operatorname{\mathsf{VAR}}$ transformer. Finally, in Section B.7, we conduct the error analysis of the Fast $\operatorname{\mathsf{VAR}}$ transformer.
B.1 Lipschitz of Polynomial